Hope you can help! I have three questions. 1. Find the slope and the y-intercept of the line. Write fractional answer in the lowest terms. 4y+4x+9=12+4x Slope is__. The y-intercept is (0,_). second question. Find the slope if it exist, of the line containing the pair of points (8,10) and (9,-8) the slope m=__. The third question. Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b (-7,8);2x=7y+5 Simplify answer use integer or fraction for any expression.
Answer by Alan3354(69443)

1. Find the slope and the y-intercept of the line. Write fractional answer in the lowest terms. 4y+4x+9=12+4x Slope is__. The y-intercept is (0,_).
4y+4x+9=12+4x
4y = 3
y = 0x + 3/4 (slope-intercept form)
The slope is zero, it's parallel to the x-axis, and the y-intercept is 3/4
-------------------------------
second question. Find the slope if it exist, of the line containing the pair of points (8,10) and (9,-8) the slope m=__.
slope m = (diff in y)/diff in x) = (-8-10)/(9-8)
= -18
-----------
The third question. Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b (-7,8);2x=7y+5 Simplify answer use integer or fraction for any expression.
2x=7y+5 Put into slope-intercept form
7y = 2x-5
The slope of the line is 2, so any line parallel will have the same slope.
Then,
y-y1 = m*(x-x1) where (x1,y1) is the point (-7,8)
y-8 = 2*(x - (-7))
y-8 = 2x + 14
y = 2x + 22